1.) Discuss the difference between the straight-line method of depreciation and the accelerated methods. Why do companies use different depreciation methods for tax reporting and financial reporting?

2.) What is the purpose of listing the account “Commitments and contingencies” on the balance sheet even through no dollar amounts appear?

3.) How is it possible for a company with positive retained earnings to be unable to pay a cash dividend?

4.) The King Corporation has total annual revenue of $800,000; expenses other than depreciation of $350,000; depreciation expense of $200,000 for tax purposes; and deprecation expense of $130,000 for reporting purposes. The tax rate is 34%. Calculate net income for reporting purposes and for tax purposes. What is the deferred tax liability?
Answer #1

1.) Discuss the difference between the straight-line method of depreciation and the accelerated methods. Why do companies use different depreciation methods for tax reporting and financial reporting?

The straight line method assumes that asset has a finite life and its ability to generate revenue is constant or same over its life. Hence, to match the expenses, the depreciation amount is equal every year over the life of the asset. Thus straight line method of depreciation results into equal and uniform depreciation over the life of the asset.

Accelerated methods assumes an asset, when new has ability to generate higher revenues and its ability to generate revenue declines with its age. Hence, as per matching concept, the depreciation expenses should be higher in the earlier years when the asset is new and lower in the later years as asset become old. Thus accelerated methods of depreciation has higher depreciation recorded in earlier years which declines every year.

Companies use different depreciation methods for tax reporting and financial reporting because:

  1. At times companies want to show a higher profits on the books hence they choose a straight line method
  2. At times companies want to defer their tax payments, so they choose accelerated depreciation methods to reduce taxable income.

-----------------------

2.) What is the purpose of listing the account “Commitments and contingencies” on the balance sheet even through no dollar amounts appear?

It's a prudent reporting practice and intends to make readers aware of the exact and complete picture of the financial health of the company while they make decisions based on the financial statements of the company. Commitments refer to crystallized but future payments to be made by the company, say for example: lease payments commitments, debt repayments in future. Contingencies are neither crystallized nor due deterministically. But we know they can translate into and get crystallized as liabilities on occurence of certain events. They are similar to contingent liabilities. Examples can be: warranty, litigations, claims from tax departments pending decision, corporate guarantees issued by the firm, letter of credits taken from banks etc.

-----------------------

3.) How is it possible for a company with positive retained earnings to be unable to pay a cash dividend?

Dividend payments are discretionary. Companies may have positive retained earnings but still decide not to pay dividends because:

  1. They may not actually have the cash to pay the dividends. This can happen if majority of revenues are yet to be collected.
  2. Companies may be preserving its internal accruals for investing into any future project or capacity expansion.

----------------------------

4.) The King Corporation has total annual revenue of $800,000; expenses other than depreciation of $350,000; depreciation expense of $200,000 for tax purposes; and depreciation expense of $130,000 for reporting purposes. The tax rate is 34%. Calculate net income for reporting purposes and for tax purposes. What is the deferred tax liability?

Net income for reporting purposes = (Revenue - expenses other than depreciation - depreciation for reporting purpose ) x (1 - Tax rate) = (800,000 - 350,000 - 130,000) x (1 - 34%) = $ 211,200

Net income for tax purposes = (Revenue - expenses other than depreciation - depreciation for tax purpose) x (1 - Tax rate) = (800,000 - 350,000 - 200,000) x (1 - 34%) = $ 165,000

Deferred tax liability = (Depreciation for tax purpose - Depreciation for reporting purpose) x tax rate = (200,000 - 130,000) x 34% = 70,000 x 34% = $ 23,800